# Frontier AI’s Impact on the Cybersecurity Landscape

### v2

@@ 5.1 Equivalence-Class-Based Prediction
Table IV shows equivalence classes across the defense lifecycle in Figure 4. Techniques developed for penetration testing, if obtained by attackers, enable more targeted attacks during reconnaissance and weaponization. Since penetration testing methods are system-specific, they provide attackers with customized attack tools. Attackers can also repurpose vulnerability detection tools to identify system weaknesses. As AI models develop stronger memorization and reasoning capabilities, they can potentially identify vulnerabilities with limited information, lowering attack thresholds for attackers with restricted target access.
@@
Beyond defense capabilities, frontier AI’s general utilities serve malicious purposes. As discussed in Section 3.2, multimodal capabilities of frontier models enable the creation of sophisticated deepfakes.
@@ 5.2 Asymmetry-Based Prediction
Asymmetry in remediation deployment and required resources. As discussed in Section 3.1, deploying remediation is time-intensive even when effective. The process requires testing for correctness, resolving dependency conflicts, global deployment, and post-deployment verification. This labor-intensive procedure demands significant resources. Legacy systems take even longer to patch, with many remaining perpetually vulnerable. Conversely, attackers can operate with minimal resource requirements. They can monitor vulnerability patches, generate exploits using public information about legacy systems, and target unpatched systems. They exploit both permanently unpatched systems and the deployment time gap between patch development and complete rollout.
@@
Different priorities of scalability and reliability. Attackers prioritize scalability, while defenders emphasize reliability. This makes AI a natural choice for attackers, as its primary capability is reducing human involvement and increasing automation. AI lowers barriers for attackers by decreasing requirements for deep expertise, specific target system knowledge, and execution time and effort. Even if an attack fails against one system, it may succeed against another. For example, in large-scale attacks, attackers may simultaneously target thousands or millions of systems (e.g., Mirai Botnet [58]). Conversely, frontier AI models face reliability limitations in robustness, transparency, and generalizability. When used for malware detection, users cannot understand why specific software is flagged as malicious or predict when false predictions will occur. This reliability deficit hinders defenders from fully trusting and deploying AI techniques in critical defense systems.
@@
This asymmetry may be magnified in critical physical infrastructure, where attackers exploit these infrastructures and propagate the damage to other systems. Beyond the asymmetry that favors attackers, certain forces and security paradigms can help shift this imbalance. For instance, large organizations typically command far greater computing resources and security expertise than individual attackers. By leveraging centralized, scalable infrastructure, they can implement comprehensive, multi-layered defenses as part of a defense-in-depth strategy. Additionally, secure-by-construction paradigms and robust defense mechanisms—such as memory safety, control flow integrity enforcement, and privilege separation—offer strong guarantees against specific vulnerability types, substantially limiting attackers’ capabilities. While these approaches help reduce attackers’ advantages, fully realizing them will likely take time.
5.3 Attack-Defense Dynamics From the Economic Perspective
Economic rationality drives both attackers and defenders—attackers strike only when expected profits exceed costs, while defenders implement protections only when benefits outweigh costs. As we have seen in Sections 5.1 and 5.2, frontier AI is changing these economic considerations significantly. By enabling more frequent and stealthy attacks, AI increases defensive costs while potentially rendering traditional defenses obsolete. This cost imbalance may force resource-constrained defenders to abandon certain protective measures and attackers will achieve higher investment returns due to diminished defenses. Appendix .2 shows a formal mathematical analysis of economic dynamics.
@@ 5.4 Evolutionary Attack-Defense Dynamics
The diverse resource and capability spectrum across attackers and defenders shapes how frontier AI impacts attack-defense dynamics over time. Attackers range from individual actors with limited resources to sophisticated state-sponsored groups with extensive funding. Similarly, defenders span from small organizations with basic security measures to large enterprises deploying advanced defense systems. Frontier AI lowers barriers to successful attacks by reducing resource and expertise requirements, likely increasing attack frequency and scale. In the early phase, well-resourced attackers amplify their capabilities by developing specialized AI models and agents, enabling more sophisticated, stealthy attacks with large-scale societal impact.
Resource-constrained defenders see limited short-term benefits from AI adoption. As detailed in Section 4.1, remediation and deployment remain time-intensive, laborious processes. Even with AI enhancing attack detection, overall defense effectiveness is constrained by these operational limitations; resource-constrained defenders struggle to leverage frontier AI to address these operational issues due to limited resources and expertise. Well-resourced defenders can leverage frontier AI to address fundamental challenges, but this requires more time. These factors indicate attackers gain advantages in the early evolutionary cycle.
Moreover, AI’s inherent randomness and the bug-finding process introduce another factor affecting dynamics. One hypothesis suggests that due to AI outputs’ statistical nature and exploration randomness, defenders and attackers may discover different vulnerabilities when analyzing the same systems. This discovery divergence creates a critical advantage for attackers, who can exploit vulnerabilities they identify while defenders address an entirely different set, even when investing more resources in bug finding. This misalignment could significantly favor attackers.
@@
Long-term, however, the evolutionary trajectory may gradually shift toward defenders as several counterfactors emerge. First, defenders could gain priority access to advanced techniques (e.g., OpenAI’s GPT-4o early access for safety researchers). As frontier AI advances, defenders benefit from improved automation, transitioning from manual patch development to AI-generated patches with automated verification. Second, as defenders use AI to systematically discover and patch vulnerabilities, systems become increasingly robust. In particular, by leveraging AI for formal verification and adopting a secure-by-design approach, defenders can build fundamentally more secure systems with provable guarantees. This continuous security improvement creates an environment where finding new vulnerabilities becomes significantly more challenging and costly for attackers. Once defenders achieve stronger system security, attackers must invest substantially more resources to find attack vectors, potentially making many attacks impractical or cost-prohibitive. Lastly, as regulatory frameworks and law enforcement mechanisms mature, attackers can face more severe penalties once caught, substantially raising failure costs and potentially deterring malicious activities.
Frontier AI will benefit attackers more than defenders in the short term. Reason 1: Frontier AI functions as a double-edged sword in cybersecurity — defensive capabilities can be repurposed for attacks (see Table IV). Reason 2: Attacker-defender asymmetry favors early AI adoption by attackers. Attackers need only succeed once, while defenders must guard against multiple threats with lower failure tolerance.
Economic perspective: Frontier AI shifts cybersecurity economics by reducing attack costs while increasing effective defense expenses, potentially forcing resource-limited defenders to abandon certain protections, though this dynamic may eventually reverse in defenders’ favor in the long term. Evolutionary dynamics: Multiple factors influence attacker-defender dynamics, including resource/capability spectrums, AI output variability, and AI development speed. Initially favoring attackers, this balance may gradually shift toward defenders over time.
6 Hybrid Systems and Security Challenges
AI agents & hybrid systems. As mentioned in Section 1, AI agents that combine LLMs with non-ML system components can enhance LLM-generated content quality and automate complex tasks. A simple example is Retrieval-Augmented Generation (RAG), which is also adopted in security applications [227]. They allow LLMs to reference external information and reduce hallucinations [167]. Beyond RAG, recent works built web agents [75, 336] for human-web interactions, coding agents [65, 312] for code writing and debugging, and personal assistants [242, 210] for daily tasks like calendar management and email. Recently, OpenAI developed deep research, an agent for conducting research tasks [209]. The security community has begun constructing multi-agent systems for complex attacks and defenses [235, 304]. An example design is to use a planning agent to decompose tasks into sub-tasks, with specialized agents collaborating with non-ML components to complete each sub-task. These examples show a shift toward increased AI integration into traditional software.
Marginal risks of hybrid systems. In Section 3.1, we identify two types of frontier AI marginal risks: augmenting existing cyber attacks and introducing new risks to hybrid systems. Figure 5 shows our taxonomy of new risks in hybrid systems, which fall into two categories: risks in foundation models and risks in AI agents/hybrid systems.
First, foundation models are vulnerable to both training-phase and testing-phase attacks. Based on existing literature [287, 267, 205, 41, 40], we outline several predominant risks that have been demonstrated in real-world commercial models. We also summarize existing automated red-teaming methods, particularly for testing-phase risks. In real-world practices, the predominant approach is to combine automated red-teaming with human red-teaming [272, 2, 276]. Attack prompts are fed to the models either directly or through (in)direct prompt injection, where attack prompts are appended after benign user prompts [183, 181].
Second, for hybrid systems, risks become more complex due to interactions between AI and symbolic components, which can be broadly classified into two categories: attack the AI component through other components and attack traditional symbolic components through AI components. Indirect prompt injection is a typical example of the first type, where attack prompts are embedded in other components rather than directly input into the model [69, 107, 341]. An example of the second type is an attack against an agent with a code generation model, a Python interpreter, a web server, and a database, designed to automatically handle data queries with proper access control. Here, an attacker might prompt the model to generate a privilege escalation code, granting illegal access to sensitive information [299]. Similarly, when an AI-driven chatbot is used in a customer service system, attackers can potentially trigger the chatbot to output inappropriate content that leads to a wrong refund [37]. Due to their complexity, solely crafting attack prompts is insufficient for effective red-teaming for hybrid systems. Recent work has started designing AI agents for red-teaming hybrid systems, mainly web agents [297, 296, 175], but this is still in its early stages. In the future, we believe there will be more automated red-teaming that works together with human-based red-teaming.
Figure 5: Taxonomy and red-teaming approaches of AI-augmented hybrid systems’ marginal risks.
Defenses for hybrid systems. Most of the existing defenses target models themselves with a focus on defending against jailbreaking attacks and prompt injection attacks. Technically speaking, such methods either conduct adversarial retraining of the target model [319, 51] or construct additional guardrail components for input or output filtering [323, 143, 182, 152]. There are only a few early explorations of defenses for hybrid systems [303, 298, 216, 339]. These methods primarily defend against attacks during tool calling, either by proactively adding data or control flow constraints[303, 298], detecting and blocking malicious tool calls [339], or implementing post-facto validation with undo features and damage confinement[216]. As hybrid systems become increasingly popular and complex, more advanced and comprehensive defenses will be in great need.
There is no work on designing defenses against hybrid systems with provable guarantees. It is a challenging task for the following reasons. First, existing formal verifications are designed for traditional symbolic programs. This makes them difficult to apply to non-symbolic programs, such as deep neural networks, which lack precisely specified properties and goals. For example, it is difficult to formally define the pre- and post-conditions for a real-world self-driving controlling system. While some early explorations exist, they lack stability and can only verify small models [154, 291]. Second, numerous studies have developed certified defenses for deep learning models [59, 165], and recent work has extended these approaches for certified defenses against jailbreaking attacks on LLMs [160, 233]. However, these efforts are still in their early stages and do not yet provide certified defenses against AI agents. Given their significance, further research is crucial to developing provable guarantees for hybrid systems.
Key takeaways. Hybrid systems: Frontier AI will drive the emergence of hybrid systems that integrate symbolic components and non-symbolic AI components. Risks: Frontier AI will introduce new marginal risks to hybrid systems at the model and system level (specified in Figure 5). Defense: Existing defenses for hybrid systems mainly focus on model-level, comprehensive system-level defenses are still lacking but will become increasingly necessary and important. None of the existing provable defenses, whether for traditional symbolic systems or machine learning, can be directly applied to hybrid systems.
TABLE V: Summary of our call to action to the security community and AI developers for improving and leveraging frontier AI to strengthen cybersecurity.
@@ 6 Hybrid Systems and Security Challenges
  Marginal risk assessment Benchmark risks in existing attacks
@@
  Lack benchmarks for fine-grained attack steps that evaluate whether AI can generate executable attacks beyond texts
@@
  Lack of attack benchmarks with carefully designed prompts for LLMs and comprehensive tool sets
Call-to-
actions
  Build benchmarks for fine-grained attack steps that cover rich target systems and possible attack paths and methods
  Combine benchmarks for individual steps to test whether AI can generate end-to-end attacks
  Involve necessary human efforts to ensure the environment correctness
  Ensure the data quality, e.g., design prompts that can possibly bypass the safety alignments of LLMs
  Design comprehensive evaluation metrics that cover both static and dynamic metrics
  Include non-ML tools that are utilized by attackers to support evaluating agent-based attacks
  Dynamically updates the benchmarks to include new attacks and support new AI tools
Benchmark new risks in hybrid systems
@@
  Lack risk categorizations and benchmarks for hybrid systems
  Lack automated red-teaming methods for hybrid systems
Call-to-
actions
  Evaluate existing model-based risk benchmarks on reasoning and small models
  Design specific benchmarks for reasoning and small models if needed
  Categorize risks for different hybrid systems and build benchmarks accordingly
  Design agent-based red-teaming methods for hybrid systems under realistic threat models
  Enhance cyber defenses Improve proactive testing
@@
  AI-driven penetration testings are far from comprehensive
  Program analysis-based detections still suffer key limitations in soundness, precisions, or scalability
@@
  Existing benchmarks lack comprehensiveness in penetration testing and have noisy data
Call-to-
actions
  Build penetration testing benchmarks with more comprehensive targets and tool supports
  Design penetration testing agents with stronger planning under comprehensive tool sets
@@
Enhance AI-based attack detection
@@
  Detect unseen attacks are still challenging
@@
  Design customized learning methods to address security-specific challenges, such as attack shifts
@@
Automate triage & remediation
@@
  Lack automation in triage and remediation deployment
  Automated patching lacks scalability and correctness
@@
Call-to-
actions
  Build multi-agent systems for root cause analysis with complex planning and program analysis tools
  Build patching agents with more accurate and comprehensive validation through differential fuzzing
  Build patching agents with formal security and correctness guarantees through symbolic executions and model checking
  Enable the agents to digest feedback from program analysis tools and iteratively refine its outcomes
@@
  Fine-tune specialized small models to replace the planning LLMs in triage and patching agent
  Leverage AI for automated deployment, e.g., testing case generation and fixing conflicts
@@
Automate provable defenses
@@
  Formal verifications (FV) are labor intensive and lack scalability
Call-to-
actions
@@
  Build secure hybrid systems Design novel security mechanisms
@@
  Lack real-time attack detection and monitoring for hybrid systems
  Lack proper system protection mechanisms, traditional protections cannot be directly used to hybrid systems
  Lack defenses with provable guarantees for hybrid systems
Call-to-
actions
@@
  Develop system- and component-level security protections for hybrid systems (e.g., generalize privilege isolation)
  Build efficient verification for AI components in the hybrid systems and the entire system as a whole
  Model developer & users Model testing & transparency
@@
Protect humans & user
@@
7 Call to Action
In this section, we provide concrete recommendations for steering frontier AI toward strengthening cybersecurity. We focus on five priorities: risk assessment as the critical first step in developing effective countermeasures, leveraging frontier AI to strengthen defenses, designing secure hybrid systems, enhancing AI model development practice and transparency, and mitigating human-related risks. Table V summarizes our recommendations.
@@ 7.1 Continuous Marginal Risk Assessment
Our first call to action is for the security community to collaborate in constructing risk assessment benchmarks. This requires covering the fine-grained risk categories in Figures 2 and 5 to enable a comprehensive assessment of AI’s marginal risks. It also demands continuous tracking of real-world AI-powered attacks and benchmark updates to reflect evolving risks. Below, we provide concrete recommendations for constructing benchmarks to assess risks of existing cyberattacks and emerging hybrid system risks.
Build systematic benchmarks incorporating fine-grained attack steps. As discussed in Section 3.1, existing attack benchmarks have deficiencies in risk coverage, data quality, and metrics. To improve risk coverage, one can leverage existing cyber ranges like MITRE’s Caldera [61] and IBM Cyber Range [141], which provide environments for various attack steps, particularly steps 3$\sim$7 where benchmarks are lacking. We can adapt these human-oriented environments with AI interfaces, tool calls, and evaluation metrics, while dynamically adding more fine-grained attack steps. Constructing attack benchmarks across diverse target systems, including private networks with different endpoints, is also essential.
For challenging tasks like CTFs, we need to incorporate tool calling capabilities and support agent construction frameworks (e.g., LangGraph [162]) to evaluate different agent designs and identify effective tools. Finally, we need to integrate existing non-AI attacks and defenses to quantify AI-specific risks, measuring both fully automated attacks and human uplift—how AI enhances human attack capabilities.
Build benchmarks and red-teaming approaches for new risks in hybrid systems. As discussed in Section 6, existing benchmarks on AI safety primarily focus on model-level risks. With the emergence of small and reasoning models, testing and extending current model-level risk benchmarks to these new models is valuable. More importantly, the community lacks risk classification and fine-grained benchmarks for hybrid systems. We must conduct fine-grained risk classification for hybrid systems and build specialized benchmarks for each category.
Since hybrid systems are highly specialized, different types will present unique risks. We recommend first categorizing hybrid systems (e.g., web agents, coding agents, personal assistants as described in Section 6) and identifying their associated risks under realistic threat models. As discussed in Section 6, we can classify risks based on their targets: AI components or symbolic components. Currently, the predominant risk is through indirect prompt injection, attacking AI through symbolic components (tools). However, there were some early attacks targeting symbolic components through AI, such as AI-generated exploits against coding agents. This underscores the critical need for thorough risk evaluations of hybrid models.
Existing red-teaming for hybrid systems only explored limited risks of specific agents (e.g., interpreter abuse of code agents [286] and prompt injection of web agents [175]). Developing automated red-teaming frameworks for broader hybrid system risks is necessary, such as assessing multi-agent system risks through game theory or designing new attacks for systems with multi-modal agents.
Continuously ensure quality and realism of benchmarks. Creating effective cybersecurity benchmarks requires careful design since agents are tasked with exploiting vulnerabilities. When benchmarks contain flaws, agents can "cheat" by completing the cyberattack task through unintended pathways. Given the complexity of cybersecurity tasks involving various files and servers, ensuring benchmarks are properly built and solvable is crucial. For instance, Transluce noted that several of the Intercode-CTF benchmark tasks are broken or unsolvable [194, 313].
Moreover, human expertise remains essential for creating quality environments and data. Proper execution environments ensure AI explores legitimate attacks rather than exploiting environment flaws (as when OpenAI-o1 hacked a CTF’s Docker container [146]). Research shows that the automated collection of vulnerable code commits without human oversight often yields non-vulnerable or security-unrelated samples [277]. Balancing quality with scale requires combining expert-created seed samples with automated methods [316]. For LLM-prompt benchmarks, one can include jailbreaking prompts to test worst-case performance. To ensure robust assessment quality, a combination of LLM judges, dynamic evaluation methods, and static program analysis techniques is recommended.
Benchmarks should reflect realistic scenarios. Real-world testing would be ideal, though this poses many challenges, including verification challenges and legal/ethical constraints. Efforts could focus on replicating real-world production systems with verification mechanisms to assess frontier AI capabilities. Complementary proxy benchmarks can also be valuable along with designing appropriate metrics—such as Cybench’s “first solve time” which compares agent performance to human expert solving times [329]. Finally, cybersecurity benchmarks require continuous updates due to its rapidly evolving threat landscape. Unlike relatively static domains like mathematics or history, cybersecurity threats transform quickly—agentic AI hacking, once considered science fiction, is now a significant concern. Therefore, even if tasks remain hidden and are not trained on, we must continuously invest in benchmarks relevant to emerging threats.

### v3

@@ 6.1 Equivalence-Class-Based Prediction
Table 3 shows equivalence classes across the defense lifecycle. Techniques developed for penetration testing, if obtained by attackers, enable more targeted attacks during reconnaissance and weaponization. Since penetration testing methods are system-specific, they provide attackers with customized attack tools. Attackers can also repurpose vulnerability detection tools to identify system weaknesses. As AI models develop stronger memorization and reasoning capabilities, they can potentially identify vulnerabilities with limited information, lowering attack thresholds for attackers with restricted target access.
@@
Beyond defense capabilities, frontier AI’s general abilities can serve malicious purposes. As discussed in Section 4.2, multimodal capabilities of frontier models enable the creation of sophisticated deepfakes.
@@ 6.2 Asymmetry-Based Prediction
Asymmetry in remediation deployment and required resources. As discussed in Section 4, deploying remediation is time-intensive even when effective. The process requires testing for correctness, resolving dependency conflicts, global deployment, and post-deployment verification. This labor-intensive procedure demands significant resources. Legacy systems take even longer to patch, with many remaining perpetually vulnerable. Conversely, attackers can operate with minimal resource requirements. They can monitor vulnerability patches, generate exploits using public information about legacy systems, and target unpatched systems. They exploit both permanently unpatched systems and the deployment time gap between patch development and complete rollout.
@@
Different priorities of scalability and reliability. Attackers prioritize scalability, while defenders emphasize reliability. This makes AI a natural choice for attackers, as its primary capability is reducing human involvement and increasing automation. AI lowers barriers for attackers by decreasing requirements for deep expertise, specific target system knowledge, and execution time and effort. Even if an attack fails against one system, it may succeed against another. For example, in large-scale attacks, attackers may simultaneously target thousands or millions of systems (e.g., Mirai Botnet [44]). Conversely, frontier AI models face reliability limitations in robustness, transparency, and generalizability [279, 222, 86]. When used for malware detection, users cannot understand why specific software is flagged as malicious or predict when false predictions will occur. This reliability deficit hinders defenders from fully trusting and deploying AI techniques in critical defense systems including decision-making processes. This asymmetry may be magnified in critical physical infrastructure, where attackers exploit these infrastructures and propagate the damage to other systems.
@@
Beyond the asymmetry that favors attackers, certain forces and security paradigms can help shift this imbalance. For instance, large organizations typically command far greater computing resources and security expertise than individual attackers. By leveraging centralized, scalable infrastructure, they can implement comprehensive, multi-layered defenses as part of a defense-in-depth strategy. Additionally, secure-by-construction paradigms and robust defense mechanisms—such as memory safety, control flow integrity enforcement, and privilege separation—offer strong guarantees against specific vulnerability types, substantially limiting attackers’ capabilities. While these approaches help reduce attackers’ advantages, fully realizing them will likely take time.
@@ 6.3 Evolutionary Attack-Defense Dynamics
The diverse resource and capability spectrum across attackers and defenders shapes how frontier AI impacts attack-defense dynamics over time. Attackers range from individual actors with limited resources to sophisticated state-sponsored groups with extensive funding. Similarly, defenders range from small organizations with basic security measures to large enterprises deploying advanced defense systems. Frontier AI lowers barriers to successful attacks by reducing resource and expertise requirements, likely increasing attack frequency and scale. In the early phase, well-resourced attackers amplify their capabilities by developing specialized AI models and agents, enabling more sophisticated, stealthy attacks with large-scale societal impact.
Frontier AI will also significantly benefit defenders. However, as detailed in Section 4.3, remediation and deployment are likely to remain time-intensive and laborious processes in the short term. Even with AI enhancing attack detection, overall defense effectiveness may be constrained by these operational limitations; resource-constrained defenders struggle to leverage frontier AI to address these operational issues due to limited resources and expertise. Well-resourced defenders can leverage frontier AI to address fundamental challenges, but this requires more time. These factors indicate attackers gain advantages in the early evolutionary cycle.
Moreover, AI’s inherent randomness and the bug-finding process introduce another factor affecting dynamics. One hypothesis suggests that due to AI outputs’ statistical nature and exploration randomness, defenders and attackers may discover different vulnerabilities when analyzing the same systems. This discovery divergence creates a critical advantage for attackers, who can exploit vulnerabilities they identify while defenders address an entirely different set. This misalignment could significantly favor attackers.
@@
Long-term, however, the evolutionary trajectory may gradually shift toward defenders as several counterfactors emerge. First, defenders could gain priority access to advanced techniques (e.g., OpenAI’s early access for safety researchers [170]). As frontier AI advances, defenders benefit from improved automation, transitioning from manual patch development to AI-generated patches with automated verification. Second, as defenders use AI to systematically discover and patch vulnerabilities, systems become increasingly robust. In particular, by leveraging AI for formal verification and adopting a secure-by-design approach, defenders can build fundamentally more secure systems with provable guarantees. This continuous security improvement creates an environment where finding new vulnerabilities becomes significantly more challenging and costly for attackers. Once defenders achieve stronger system security, attackers must invest substantially more resources to find attack vectors, potentially making many attacks impractical or cost-prohibitive. Lastly, as regulatory frameworks and law enforcement mechanisms mature, attackers can face more severe penalties once caught, substantially raising failure costs and potentially deterring malicious activities.
Table 4: Summary of our calls to action. We bold several key and immediate recommendations.
@@
Marginal
risk
assessment
Benchmark
risks in
existing
attacks
@@
  Lack benchmarks for comprehensive attack steps that evaluate whether AI can generate executable attacks
@@
  Lack of dynamic environments that support agent execution and calling tools
Calls-to-
action
  Build comprehensive benchmarks, covering fine-grained attack steps and various systems
  Involve necessary human effort to ensure environment correctness and data quality
  Design end-to-end dynamic environments to support agentic execution; providing non-ML tools and agent scaffolds
  Periodically update benchmarks to include new attacks and support new AI tools
Benchmark
new risks
in AI agents
@@
  Lack risk categorizations, benchmarks, and automated red-teaming methods for AI agents
Calls-to-
action
  Categorize risks for different AI agents and build benchmarks accordingly
  Design agent-based red-teaming methods for advanced models and AI agents under realistic threat models
Enhance
cyber
defenses
Improve
proactive
testing
@@
  AI-driven penetration testing is still at an early stage
  Program analysis-based detections still suffer a trade-off in soundness, precision, or scalability
@@
  Existing benchmarks lack comprehensiveness in penetration testing and have noisy data for vulnerability detection
Calls-to-
action
  Build comprehensive penetration testing benchmarks, covering more tasks and systems
  Design penetration testing agents with domain-specific tool sets
@@
Enhance
AI-based
attack
detection
@@
  Detect unseen attacks or long-tail tasks with fewer data is still challenging
@@
  Design customized learning methods to address security-specific challenges, such as attack shifts, long tail problems
@@
Automate
triage &
remediation
@@
  Lack automation in triage and remediation deployment; patching agents are simple and lack correctness
@@
Calls-to-
action
  Build multi-agent systems for triage and patching with program analysis tools
  Build patching agents with validation through differential fuzzing, symbolic executions, and model checking
  Enable the agents to digest feedback from program analysis tools and iteratively refine their outcomes
@@
  Fine-tune specialized reasoning models under agentic systems with SFT or RL
  Leverage AI for automated deployment and operation, e.g., testing case generation and fixing conflicts
@@
Automate
provable
defenses
@@
  Formal verifications (FV) are labor-intensive and lack scalability
Calls-to-
action
@@
Build
secure
AI
agents
Design
novel
security
mechanisms
@@
  Lack real-time attack detection and monitoring for AI agents
  Lack proper system protection mechanisms, traditional protections cannot be directly used to AI agents
  Lack defenses with provable guarantees for AI agents
Calls-to-
action
@@
  Develop system- and component-level security protections for AI agents (e.g., generalize privilege isolation)
  Build efficient verification for AI components in the AI agents and the entire system as a whole
Model
developer
& users
Model
testing &
transparency
@@
Protect humans &
user
@@
7 Calls to Action
We provide concrete recommendations for steering frontier AI toward strengthening cybersecurity from five priorities: risk assessment, leveraging frontier AI to strengthen defenses, designing secure AI agent systems, enhancing AI model development practice and transparency, and mitigating human-related risks. Table 4 summarizes our recommendations.
@@ 7.1 Continuous Marginal Risk Assessment
We recommend that the security and AI communities collaborate on constructing comprehensive risk assessment benchmarks. These should cover the fine-grained risk categories in Figure 2 and diverse systems, include end-to-end attack scenarios with executable environments where agents can take sequential actions, and be updated periodically to track new attacks.
[…]
Build benchmarks and red-teaming approaches for new risks in AI agents. We need to conduct fine-grained risk classification for agents and build specialized benchmarks for each category, given that agents are highly specialized (e.g., web agents, coding agents, and personal assistants). We can also classify risks based on their targets: AI components or symbolic components. Currently, the predominant risk is through indirect prompt injection, attacking AI through symbolic components (tools). However, there are only a few early attacks targeting symbolic components through AI, such as AI-generated exploits against coding agents. This underscores the critical need for thorough risk evaluations of agents. Existing red-teaming for agents only explores limited risks of specific agents (e.g., interpreter abuse of code agents [236], prompt injection of web agents [137, 69], indirect prompt injection [243], and taint-style vulnerability [143] of various agents). Developing automated red-teaming frameworks for broader agentic risks is necessary, such as assessing multi-agent system risks through game theory or designing new attacks for systems with multi-modal agents.
Continuously ensure quality and realism of benchmarks. If a security benchmark is flawed, agents can “cheat” by completing the task via unintended pathways rather than the intended exploit. Given the complexity of these tasks—often spanning multiple files, services, and hosts—benchmarks must be carefully constructed, validated, and demonstrably solvable. For instance, Transluce noted that several of the Intercode-CTF benchmark tasks are broken or unsolvable [154, 257]. Moreover, human expertise remains essential for creating quality environments and data. Research shows that the automated collection of vulnerable code commits without human oversight often yields non-vulnerable or security-unrelated samples [228]. Balancing quality with scale requires combining expert-created seed samples with automated methods [260]. To ensure robust assessment quality, a combination of LLM judges, dynamic evaluation methods, and static program analysis techniques is recommended. Benchmarks should reflect realistic scenarios; real-world testing would be ideal, but raises challenges like verification difficulties and legal/ethical constraints. A practical approach is to replicate production systems in controlled environments with built-in verification mechanisms to assess frontier AI capabilities. Finally, cybersecurity benchmarks must be continuously updated to keep pace with the rapidly evolving threat landscape. Timely updates surface emerging AI-enabled capabilities, allowing practitioners to prioritize mitigations and validate defenses before widespread exploitation.